# Monadic Second Order Logic of the Real Order with the Set Quantifier ranging over the Borel Sets

### v2

@@ 1. Introduction
Monadic Second Order logic (MSO) is defined like the first order logic (FO), but with an addition of set variables and the additional atomic formulas of the form $x\in{\mathbf{A}}$ and $(\exists{\mathbf{A}})\psi({\mathbf{A}})$ for a set-quantifier-free $\psi({\mathbf{A}})$. This logic is a fragment of the Second Order Logic (SO) that has proven sufficiently rich to express various useful mathematical statements. The results of this paper show, for example, that for any given $\alpha<\omega_{1}$ the Determinacy of the $\alpha$-level of the Borel hierarchy is expressible as a sentence in the MSO theory S2S of the Cantor tree with the two successor functions (Corollary 6.3). On the other hand, MSO is close enough to FO to have many properties that let us work with the logic in practice, in particular many instances of decidability. The decidability of various logics is a question stemming directly from Hilbert’s programme, most famously through Hilbert’s 10th Problem and Yuri Matijashevich’s 1970 solution to it [16]. The full SO is most often undecidable. In the case of SO arithmetic, so the SO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, the undecidability was proven by Alonzo Church in [4]. In contrast, the MSO theory of various structures is decidable, including the MSO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, which was famously proved decidable by J. Richard Büchi in [1]. In his proof Büchi introduced an important method, which is to associate to an MSO theory the runs of an automaton. A posteriori, one could say that the most important property of MSO is its closeness to the theory of automata. This makes MSO a logic much studied in theoretical computer science. The survey article [7] by Yuri Gurevich presents many properties of the MSO logic.
@@
The paper is organised as follows. In §2 we define well-levelled orders. They are a class of well-founded partial orders and every well-order and every tree is well-levelled. We present the basic properties of well-levelled partial orders and show some examples, to the extent that will be useful for the main result and slightly more, as a reference. In §3 we show how to use these ideas to code Borel subsets of the reals. In §4 we show how to use this coding, combined with the ideas from the original proof by Rabin, for the main decidability result. The main results are presented in §5 in Theorem 5.1 and Corollary 5.3.
@@ 2. Well-levelled orders and their basic properties
(6) Let $T$ be a rooted tree of height $\omega$ and we define by induction on $n\geq 1$ what is $T^{n}$. Let $\Lambda$ denote the root of $T$.
@@
For higher powers, such as $\omega$, we need more notation. Hence we comment on how the projections as described in the above relate to the natural restriction maps, which will then allow us to liberate the word “projection” for other purposes in the paper.
@@ 2.1. The Polish Topology of $T^{1+\alpha}]$
For countable ordinals $\alpha$, each $T^{1+\alpha}]$ has a naturally defined complete metric, so the resulting topology, which we shall denote by $\tau_{\alpha}$ makes $T^{1+\alpha}]$ into a Polish space. To start the proof of this, recall ([9, pg. 7] that the Cantor space $[T]$ has a complete metric given by $d_{0}=d$ where,
@@ 3. Coding of the Borel Sets
To start the definition, we recall the well-known coding of the closed subsets $H$ of the Cantor set $2^{\omega}$. We identify the product $2^{\omega}$ of $\omega$ many copies of $2=\{0,1\}$ in the product topology with a topological space obtained on the set $^{\omega}2$ of $\omega$-sequences of 0s and 1s, with a basis of the topology given by the sets $s]=\{f\in{}^{\omega}2:\,s\subseteq f\}$ for $s$ a finite partial function from $\omega$ to 2 (equivalently, we can restrict our attention to the $s$ whose domain is some natural number). It is usual to denote both of these objects by $2^{\omega}$ and call them both the Cantor set.
@@
$\{t(\beta_{n}):\,n<\omega,t\in\bigcup S,t\upharpoonright(\beta_{n}\setminus\alpha)=\Lambda\}\setminus\{\Lambda\}$
@@
For $\alpha>0$ a limit, suppose that we have two $\Sigma_{\alpha}$-sets $H_{l}=\bigcup_{n<\omega}H^{l}_{n}$ for $l<2$ such that each $H^{l}$ is coded using the sequence $\langle H^{l}_{n}:\,n<\omega\rangle$. Hence, each of the sets $H_{n}^{l}$ is a $\Pi_{\beta^{l}_{n}}$-set for some $\beta^{l}_{n}<\alpha$, the sequences $\langle H^{l}_{n}:\,n<\omega\rangle$ are $\subseteq$-increasing and $\sup_{n<\omega}\beta^{l}_{n}=\alpha$. Moreover, the sequence $\langle\beta^{l}_{n}:\,n<\omega\rangle$ does not depend on $l$ and is the canonically chosen sequence used to define codes in $T^{\alpha}]$.
If both of the sets $H_{0},H_{1}$ are $\bigcup_{\beta<\alpha}(\Sigma_{\beta}\cup\Pi_{\beta})$-sets the situation is already covered by the induction hypothesis. Otherwise, at least one of them is a $\Sigma_{\alpha}$-set without being a $\bigcup_{\beta<\alpha}(\Sigma_{\beta}\cup\Pi_{\beta})$-set. We can assume by symmetry that it is $H_{0}$. If $H_{1}=\emptyset$ then the situation is trivial, so let us assume that $H_{1}\neq\emptyset$. Let $a\in H_{1}$ be any element (for definability purposes, we can choose it to be the first in the lexicographic order of $[T]$).
@@
Now suppose that neither $H_{0}$ nor $H_{1}$ are $\bigcup_{\beta<\alpha}(\Sigma_{\beta}\cup\Pi_{\beta})$-sets. We have that $H_{0}\cap H_{1}=\bigcup_{n<\omega}(H^{0}_{n}\cap H^{1}_{n})$, as in the case of $\Sigma_{\alpha+1}$-sets above. Then by the induction hypothesis applied to $\beta_{n}$, for each $n$ we can write, with the same conventions as before $c(H^{0}_{n}\cap H^{1}_{n})=c(H^{0}_{n})\cap c(H^{1}_{n})$. We consider which branches $\pi$ of $T^{\alpha}$ we use in constructing $c(H)$ in the definition of the Borel codes, and we verify that we obtain exactly $c(H_{0}\cap H_{1})=c(H_{0})\cap c(H_{1})$.
@@ 4. Decidability of $T^{1+\alpha}$
Towards the goal of answering in the affirmative the question of the decidability of the monadic second order logic on the reals where the monadic quantifier ranges over the Borel subsets of the reals, we develop translations into sentences of $S2S$. The latter we know decidable by [19]. The main translation is introduced in Definition 4.1.
@@ 4.1. The recursive correspondance between MSO formulas over $T$ and Rabin automata
It was shown in [19, pg.8] that both $\leq$ and $\leq_{\rm lex}$ definable using just $r_{0}$ and $r_{1}$. It follows that the MSO theory of the ordered Cantor tree along with the lexicographic order is a sub-theory of $S2S$. The main theorem of [19] is Theorem 1.7, which shows that to every formula of $S2S$ there is an associated Rabin automaton which represents the formula. To recall, this means that there is an elementary recursive procedure which to every such formula $F({\bf A}_{1},\ldots{\bf A}_{n})$, where ${\bf A}_{1},\ldots{\bf A}_{n}$ are set variables, associates a Rabin automaton ${\mathfrak{A}}_{F}$ which accepts exactly the set of all representations of tuples $(A_{1},\ldots A_{n})$ such that ${\mathfrak{N}}_{2}\vDash F(A_{1},\ldots A_{n})$. Let us recall what is meant by the Rabin automaton and representations (see [19, pg. 6]).
@@ 4.2. Towards the automata associated to $n$-tuples of subsets of $T^{1+\alpha}$
Our main auxiliary theorem is to show that the theory of every $T^{1+\alpha}$ with the set predicates ranging of $\Pi_{\alpha}$-sets is decidable, and also the theory of $T^{\alpha}\cdot B$ with the set predicates ranging of $\Sigma_{\alpha+1}$-sets. This has to be made precise by specifying which operations are allowed, which we do below, but the idea is to do for these theories what Rabin has done for $S\omega S$.
@@ 4.3. Translating into $S2S$
    if $\alpha=\beta+1$, the projections $x\mapsto b(x)$ from $T^{1+\beta+1}\to T$ and $x\mapsto p(x)$ from $T^{1+\beta+1}\to T^{1+\beta}$ (see Notation 3.1) are MSO-definable,
@@
    suppose that $\langle F_{\beta}({\mathbf{A}}_{1},\ldots{\mathbf{A}}_{n}):\,\beta<\alpha\rangle$ is a sequence of MSO-formulas. Then the function $\beta\mapsto F_{\beta}({\mathbf{A}}_{1},\ldots{\mathbf{A}}_{n})$ is MSO-definable
@@
[…]
Coming back to our $\alpha$, we use these observations to give an MSO-representation of $(T^{\alpha},\leq_{\alpha})$ within $T$. By the induction hypothesis, we have an MSO-representation of each $(T^{\beta},\leq_{\beta})$ for $\beta<\alpha$ AND the elements of $T^{\beta}$ are represented by certain $\beta$-sequences of $T$. Let $(A_{\alpha},\prec_{\alpha})$ be an MSO-definable subset of $T$ ordered by $\prec_{\alpha}$ in order type $\alpha$. We shall define a formula $G_{\alpha}(\varrho)$:
@@
(3) Given that we have established that being in $T^{1+\beta}$ is MSO-definable, this item is similar to the classical case: $b(\bar{x})=y$ if there is $\bar{z}$ such that $\bar{x}=\bar{z}\frown y$ and $\bar{z}$ is in $T^{1+\beta}$. Similarly for the projection.
@@ 4.5. Subsets of $T^{1+\alpha}$ and branches in $T^{1+\alpha}]$
Let us first remark that for every $\alpha$ there is a formula ${\rm Path}_{1+\alpha}(\mathbf{B})$ such that for $B\subseteq T^{1+\alpha}$ we have ${\mathfrak{N}}_{2}\models(t{\rm Path}_{1+\alpha})(f_{1+\alpha}(B))$ iff $B$ is a path of $T^{1+\alpha}$. This formula is
@@
We shall use the letter $\pi$ throughout to denote a branch, specifying in which $\leq_{\alpha}$ unless clear from the context. The following observation will be useful in the proof of Theorem 5.1.
@@ 4.6. Choosing the $\sigma$-representation by open sets in an MSO-definable way
For the case $\alpha+1$, let us first consider the $\Sigma_{\alpha+1}$-sets $H_{0}$ and $H_{1}$, represented using the sequences $\langle H^{l}_{n}:\,n<\omega\rangle$ for $l<2$ of $\Pi_{\alpha}$-sets. Going back to the proof of Lemma 3.8, in the case that $H_{0}\cap H_{1}\neq\emptyset$ we have that given any codes $\langle H^{l}_{n}:\,n<\omega\rangle$ for $l<2$, we can in a definable way, by looking for the first $n$ such that $H^{0}_{n}\cap H^{1}_{n}\neq\emptyset$, change the codes so that $\langle H^{0}_{n}\cap H^{1}_{n}:\,n<\omega\rangle$ is a code for $H^{0}\cap H^{1}$. By the inductive hypothesis for the $\Pi_{\alpha}$-sets, for every $n<\omega$ we have an MSO-definable way of choosing the codes for $H^{0}_{n}$ and $H^{1}_{n}$ so that $c(H^{0}_{n}\cap H^{1}_{n})=c(H^{0}_{n})\cap c(H^{1}_{n})$, say by some MSO-definable function $F_{n}$. By Lemma 4.2(4), we can find an MSO-definable function $F$ which lets us perform the whole sequence $\langle F_{n}:\,n<\omega\rangle$ of the changes simultaneously. Then $F$ gives us the codes for $H^{0}$ and $H^{1}$ which are as required. In the case that $H_{0}\cap H_{1}=\emptyset$, we have that for each $n$ the intersection $H^{0}_{n}\cap H^{1}_{n}$ is empty, hence we apply the induction hypothesis for the $\Pi_{\alpha}$-sets in the same way as in the previous case.

### v3

@@ 1. Introduction
Monadic Second Order logic (MSO) is defined like the first order logic (FO), but with an addition of set variables and the additional atomic formulas of the form $x\in{\mathbf{A}}$ and $(\exists{\mathbf{A}})\psi({\mathbf{A}})$ for a set-quantifier-free $\psi({\mathbf{A}})$. This logic is a small fragment of the Second Order Logic (SO), yet it has proven sufficiently rich to express various useful mathematical statements. The results of this paper show, for example, that for any given $\alpha<\omega_{1}$ the Determinacy of the $\alpha$-level of the Borel hierarchy is expressible as a sentence in the MSO theory S2S of the Cantor tree with the two successor functions (Corollary 6.3). On the other hand, MSO is close enough to FO to have many properties that let us work with the logic in practice, in particular many instances of decidability. The decidability of various logics is a question stemming directly from Hilbert’s programme, most famously through Hilbert’s 10th Problem and Yuri Matijashevich’s 1970 solution to it [16]. The full SO is most often undecidable. In the case of SO arithmetic, so the SO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, the undecidability was proven by Alonzo Church in [4]. In contrast, the MSO theory of various structures is decidable, including the MSO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, which was famously proved decidable by J. Richard Büchi in [1]. In his proof Büchi introduced an important method, which is to associate to an MSO theory the runs of an automaton. A posteriori, one could say that the most important property of MSO is its closeness to the theory of automata. This makes MSO a logic much studied in theoretical computer science. The survey article [7] by Yuri Gurevich presents many properties of the MSO logic.
@@
The paper is organised as follows. In §2 we define well-levelled orders. They are a class of well-founded partial orders and every well-order and every tree is well-levelled. We present the basic properties of well-levelled partial orders and show some examples, to the extent that will be useful for the main result and slightly more, as a reference. In §3 we show how to use these ideas to code Borel subsets of the reals. In §4 we show how to use this coding, combined with the ideas from the original proof by Rabin, for the main decidability result. The main results are presented in §5 in Theorem 5.1 and Corollary 5.3. For the stronger version of the main result, which is Corollary 5.3, we had to leave S2S and use an additional result of Büchi [3], stating that the MSO theory of $(\omega_{1},<)$ is decidable.
@@ 2. Well-levelled orders and their basic properties
(6) Let $T$ be a rooted tree and we define by induction on $n\geq 1$ what is $T^{n}$. Let $\Lambda$ denote the root of $T$.
@@ 2.1. The Polish Topology of $T^{1+\alpha}]$
For countable ordinals $\alpha$, each $T^{1+\alpha}]$ has a naturally defined complete metric, so the resulting topology, which we shall denote by $\tau_{\alpha}$ makes $T^{1+\alpha}]$ into a Polish space. To start the proof of this, recall ([9, pg. 7]) that the Cantor space $[T]$ has a complete metric given by $d_{0}=d$ where,
@@ 3. Coding of the Borel Sets
To start the definition, we recall the well-known coding of the closed subsets $H$ of the Cantor set $2^{\omega}$. We identify the product $2^{\omega}$ of $\omega$ many copies of $2=\{0,1\}$ in the product topology with a topological space obtained on the set $^{\omega}2$ of $\omega$-sequences of 0s and 1s, with a basis of the topology given by the sets $s]=\{f\in{}^{\omega}2:\,s\subseteq f\}$ for $s$ a finite partial function from $\omega$ to 2 (equivalently, we can restrict our attention to the $s$ whose domain is some natural number as a sub-basis). It is usual to denote both of these objects by $2^{\omega}$ and call them both the Cantor set.
@@
$\{t(\beta_{n}):\,n<\omega,t\in\bigcup S,t\upharpoonright(\beta_{n}\setminus\alpha)\equiv\Lambda\}\setminus\{\Lambda\}$
@@
For $\alpha>0$ a limit, suppose that we have two $\Sigma_{\alpha}$-sets $H_{l}=\bigcup_{n<\omega}H^{l}_{n}$ for $l<2$ such that each $H^{l}$ is coded using the sequence $\langle H^{l}_{n}:\,n<\omega\rangle$. Hence, each of the sets $H_{n}^{l}$ is a $\Pi_{\beta^{l}_{n}}$-set for some $\beta^{l}_{n}<\alpha$, the sequences $\langle H^{l}_{n}:\,n<\omega\rangle$ are $\subseteq$-increasing and $\sup_{n<\omega}\beta^{l}_{n}=\alpha$. Moreover, the sequence $\langle\beta^{l}_{n}:\,n<\omega\rangle$ does not depend on $l$ and is the canonically chosen sequence $\langle\beta_{n}:\,n<\omega\rangle$ used to define codes in $T^{\alpha}]$.
If both of the sets $H_{0},H_{1}$ are $\bigcup_{\beta<\alpha}(\Sigma_{\beta}\cup\Pi_{\beta})$-sets the situation is already covered by the induction hypothesis. Otherwise, at least one of them is a $\Sigma_{\alpha}$-set without being a $\bigcup_{\beta<\alpha}(\Sigma_{\beta}\cup\Pi_{\beta})$-set. We can assume by symmetry that it is $H_{0}$. If $H_{1}=\emptyset$ then the situation is trivial, so let us assume that $H_{1}\neq\emptyset$. First suppose that $H_{0}\cap H_{1}\neq\emptyset$. Let $a\in H_{0}\cap H_{1}$ be any element (for definability purposes, we can choose it to be the first in the lexicographic order of $[T]$).
@@
Now suppose that neither $H_{0}$ nor $H_{1}$ are $\bigcup_{\beta<\alpha}(\Sigma_{\beta}\cup\Pi_{\beta})$-sets. We have that $H_{0}\cap H_{1}=\bigcup_{n<\omega}(H^{0}_{n}\cap H^{1}_{n})$, as in the case of $\Sigma_{\alpha+1}$-sets above. Then by the induction hypothesis applied to $\beta_{n}$, for each $n$ we can write, with the same conventions as before $c(H^{0}_{n}\cap H^{1}_{n})=c(H^{0}_{n})\cap c(H^{1}_{n})$. Similarly as in the above, we can handle the possibility of the first finitely many $n$ such that $H^{0}_{n}\cap H^{1}_{n}=\emptyset$, so let us assume that this has been done. At the end we consider which branches $\pi$ of $T^{\alpha}$ we use in constructing $c(H)$ in the definition of the Borel codes, and we verify that we obtain exactly $c(H_{0}\cap H_{1})=c(H_{0})\cap c(H_{1})$.
It remains to deal with the case $H_{0}\cap H_{1}=\emptyset$. This implies that there is $n_{0}$ such that for all $n\geq n_{0}$ we have that $H^{0}_{n}\cap H^{1}_{n}=\emptyset$ and hence by the induction hypothesis we can assume that $c(H^{0}_{n_{0}})\cap(H^{1}_{n_{0}})=\emptyset$. By the construction of codes at limit ordinals it follows that $c(H_{0})\cap(H_{1})=\emptyset$.
@@ 4. Decidability of $T^{1+\alpha}$
Towards the goal of answering in the affirmative the question of the decidability of the monadic second order logic on the reals where the monadic quantifier ranges over the Borel subsets of the reals, we develop various translations into sentences of $S2S$. The latter we know decidable by [19]. The main translation is introduced in Definition 4.1.
@@ 4.1. The recursive correspondance between MSO formulas over $T$ and Rabin automata
It was shown in [19, pg.8] that both $\leq$ and $\leq_{\rm lex}$ are MSO-definable using just $r_{0}$ and $r_{1}$. It follows that the MSO theory of the ordered Cantor tree along with the lexicographic order is a sub-theory of $S2S$. The main theorem of [19] is Theorem 1.7, which shows that to every formula of $S2S$ there is an associated Rabin automaton which represents the formula. To recall, this means that there is an elementary recursive procedure which to every such formula $F({\bf A}_{1},\ldots{\bf A}_{n})$, where ${\bf A}_{1},\ldots{\bf A}_{n}$ are set variables, associates a Rabin automaton ${\mathfrak{A}}_{F}$ which accepts exactly the set of all representations of tuples $(A_{1},\ldots A_{n})$ such that ${\mathfrak{N}}_{2}\vDash F(A_{1},\ldots A_{n})$. Let us recall what is meant by the Rabin automaton and representations (see [19, pg. 6]).
@@ 4.2. Towards the automata associated to $n$-tuples of subsets of $T^{1+\alpha}$
Our main auxiliary theorem is to show that the theory of every $T^{1+\alpha}$ with the set predicates ranging over $\Pi_{\alpha}$-sets is decidable, and also the theory of $T^{\alpha}\cdot B$ with the set predicates ranging over $\Sigma_{\alpha+1}$-sets. This has to be made precise by specifying which operations are allowed, which we do below, but the idea is to do for these theories what Rabin has done for $S\omega S$.
@@ 4.3. Translating into $S2S$
    the projections $x\mapsto b(x)$ from $T^{1+\alpha+1}\to T$ and $x\mapsto p(x)$ from $T^{1+\alpha+1}\to T^{1+\alpha}$ (see Notation 3.1) are MSO-definable,
@@
    suppose that $\langle F_{\beta}({\mathbf{A}}_{1},\ldots{\mathbf{A}}_{n}):\,\beta<\alpha\rangle$ is a sequence of MSO-formulas. Then the function $\beta\mapsto F_{\beta}({\mathbf{A}}_{1},\ldots{\mathbf{A}}_{n})$ is MSO-definable.
@@
Now we have that having the order type $\beta$ is MSO-definable over $(\beta,<)$ by the formula $F(X)=(\forall x)(x\in X)$. Therefore there is a formula $F_{\beta}(X,Y)$ such that $F_{\beta}(X,Y)$ holds in $T$ iff $X\subseteq T$, $Y$ is an order relation on $X$ and $X$ has order type $\beta$ in the relation $Y$.
Coming back to our $\alpha$, we use these observations to give an MSO-representation of $(T^{\alpha},\leq_{\alpha})$ within $T$. By the induction hypothesis, we have an MSO-representation of each $(T^{\beta},\leq_{\beta})$ for $\beta<\alpha$ AND the elements of $T^{\beta}$ are represented by certain $\beta$-sequences of $T$. Let $(A_{\alpha},\prec_{\alpha})$ be an MSO-definable subset of $T$ ordered by $\prec_{\alpha}$ in the order type $\alpha$. We shall define a formula $G_{\alpha}(\varrho)$:
@@
(3) Given that we have established that being in $T^{1+\alpha}$ is MSO-definable, this item is similar to the classical case: $b(\bar{x})=y$ if there is $\bar{z}$ such that $\bar{x}=\bar{z}\frown y$ and $\bar{z}$ is in $T^{1+\alpha}$. Similarly for the projection.
@@ 4.5. Subsets of $T^{1+\alpha}$ and branches in $T^{1+\alpha}]$
Let us first remark that for every $\alpha$ there is a formula ${\rm Path}_{1+\alpha}(\mathbf{B})$ such that for $B\subseteq T^{1+\alpha}$ we have ${\mathfrak{N}}_{2}\models(t{\rm Path}_{1+\alpha})(f{1+\alpha}(B))$ iff $B$ is a path of $T^{1+\alpha}$. This formula is
@@
From now on we shall use the letter $\pi$ to denote a branch, specifying in which $\leq_{\alpha}$ unless clear from the context. The following observation will be useful in the proof of Theorem 5.1.
@@ 4.6. Choosing the $\sigma$-representation by open sets in an MSO-definable way
For the case $\alpha+1$, let us first consider the $\Sigma_{\alpha+1}$-sets $H_{0}$ and $H_{1}$, represented using the sequences $\langle H^{l}_{n}:\,n<\omega\rangle$ for $l<2$ of $\Pi_{\alpha}$-sets. Going back to the proof of Lemma 3.8, in the case that $H_{0}\cap H_{1}\neq\emptyset$ we have that given any codes $\langle H^{l}_{n}:\,n<\omega\rangle$ for $l<2$, we can in an MSO-definable way, by looking for the first $n$ such that $H^{0}_{n}\cap H^{1}_{n}\neq\emptyset$, change the codes so that $\langle H^{0}_{n}\cap H^{1}_{n}:\,n<\omega\rangle$ is a code for $H^{0}\cap H^{1}$. By the inductive hypothesis for the $\Pi_{\alpha}$-sets, for every $n<\omega$ we have an MSO-definable way of choosing the codes for $H^{0}_{n}$ and $H^{1}_{n}$ so that $c(H^{0}_{n}\cap H^{1}_{n})=c(H^{0}_{n})\cap c(H^{1}_{n})$, say by some MSO-definable function $F_{n}$. By Lemma 4.2(4), we can find an MSO-definable function $F$ which lets us perform the whole sequence $\langle F_{n}:\,n<\omega\rangle$ of the changes simultaneously. Then $F$ gives us the codes for $H^{0}$ and $H^{1}$ which are as required. In the case that $H_{0}\cap H_{1}=\emptyset$, we have that for each $n$ the intersection $H^{0}_{n}\cap H^{1}_{n}$ is empty, hence we apply the induction hypothesis for the $\Pi_{\alpha}$-sets in the same way as in the previous case.